19) A company purchased a delivery van for $23,000 with a salvage value of $3,000. It...

19)

A company purchased a delivery van for $23,000 with a salvage value of $3,000. It has an estimated life of 5 years. Using the straight-line method, how much depreciation expense should the company recognize each year?  

$5,000

$1,000

$4,000

$1,533

20)

A company sold $20,000 of equipment, accepting a 30-day note bearing interest at 12% per year (assume a 360 day year) for the $20,000. The entry to record the receipt of the amount collected at maturity will include which of the following, assuming no interest had been accrued.?

Credit interest revenue $200

Credit interest revenue $2,400

Credit Notes Receivable $20,200

Debit Cash for $20,000
Answer #1
19) Straight Line Depreciation = Purchase Price of Asset - Approximate salvage value/Estimated Useful Life of Asset
Purchase Price of Asset 23000
Approximate salvage value 3000
Estimated Useful life of asset 5 years
Straight Line Depreciation = 23000-3000/5 years
= 23000-3000
= 20000
= 20000/5
Depreciation Expenses = 4000
Correct answer : Option C = 4000
Debit Credit
20) Account Receivable 20200
Notes Receivable 20000
Interest Revenue 200
(20000*12%*30/360 = 200)
Debit to Accounts receivable for 20200
Credit to notes Receivable for 20200
Debit to cash for 20200
Correct answer : option C =credit notes receivable 20200